# Causal Inference in Longitudinal Data under Unknown Interference

### v4

@@ Abstract
In longitudinal studies where units are embedded in space or a social network, interference may arise, meaning that a unit’s outcome can depend on treatment histories of others. The presence of interference poses significant challenges for causal inference, particularly when the interference structure—how a unit’s outcome responds to others’ influences—is complex, heterogeneous, and unknown to researchers. This paper develops a general framework for identifying and estimating both direct and spillover effects of treatment histories under minimal assumptions about the interference structure. We define a class of policy-relevant causal estimands and show that they can be represented by a modified marginal structural model (MSM). Under the standard assumption of sequential exchangeability, these estimands are identifiable and can be estimated using inverse probability weighting (IPW). We derive conditions for consistency and asymptotic normality of the estimators and provide procedures for constructing Wald-type confidence intervals with valid coverage in large samples. The method’s utility is demonstrated through applications in both social science and biomedical settings.
@@ 1 Introduction
Empirical researchers are often interested in estimating the causal effects of treatment histories in longitudinal data. For example, public health researchers may seek to determine whether a history of smoking increases the likelihood of developing cardiovascular diseases, while social scientists want to understand how past experiences with government programs shape voting behavior. Under the standard assumption of sequential exchangeability—in either experimental or observational studies—these effects can be represented by a marginal structural model (MSM) and consistently estimated through inverse probability weighting (IPW) (Robins, Hernan and Brumback, 2000; Hernán and Robins, 2010).
@@
However, the classic approach relies on the stable unit treatment value assumption (SUTVA), which rules out the possibility that the outcome of a unit is affected by treatment histories of others, a phenomenon known as spillover or interference (Cox, 1958). In contexts where units are embedded in geographic space or social networks, interference is likely present, raising questions on how researchers can describe and estimate the direct or spillover effects a treatment history generates. Most existing solutions assume a known interference structure, typically formalized as an “exposure mapping” that specifies how each unit’s outcome depends on the treatments received by a fixed set of others (Aronow and Samii, 2017; Ogburn et al., 2020). For example, in studies of second-hand smoke, it is often assumed that an individual’s health is influenced only by the average smoking behavior of their immediate contacts.
In many applications, nevertheless, the interference structure is complex, heterogeneous, and unknown to researchers. Individuals may be affected by second-hand smoke from socially or spatially distant others, depending on their unobserved daily activities. In such cases, which we refer to as “unknown interference,” exposure mappings may fail to capture the relevant sources of influence, undermining identification and estimation.
In this paper, we present a general framework to address this problem by integrating existing methods in longitudinal data analysis with recent advancements on causal inference under unknown interference (Sävje, Aronow and Hudgens, 2021; Wang et al., 2025). We introduce a novel estimand, the average marginalized response (AMR), which can be used to capture either the direct or spillover effects and remains well-defined even when the interference structure is unknown. The only requirement is the existence of a proximity metric between any two units (e.g., spatial distance, network path length, or shared group membership), which is necessary for indexing spillover effects but but imposes no restrictions on the form or complexity of the interference structure.
For a fixed proximity level $d\geq 0$, the AMR summarizes how, on average, the expected outcomes of a unit’s neighbors at that proximity level respond to the unit’s treatment history. As a function of treatment history, AMR can naturally be represented by a MSM, where linear combinations of the parameters reflect the causal effect of a unit’s treatment history on the outcomes of its neighbors at proximity level $d$. By evaluating AMRs and their associated MSMs across different values of $d$, researchers can examine how the effects of treatment histories vary over the proximity to one’s neighbors and distinguish the direct effect ($d=0$) from spillover effects ($d>0$) along both temporal and cross-unit dimensions.
@@
We then demonstrate that both the AMR and the parameters in the associated MSM can be identified under the familiar assumption of sequential exchangeability. This identification result enables estimation of MSM parameters using IPW estimators, such as weighted least squares (WLS), where we regress the average outcome of a unit’s neighbors at the proximity level of $d$ on indicators of its own treatment history, weighted by the inverse probability of that history being observed. We prove that this estimator exhibits desirable large-sample properties—specifically, consistency and asymptotic normality—when either the influence of a unit’s treatment decays sufficiently quickly over proximity or neighborhood sizes grow slowly with the sample size, using recently developed central limit theorems for dependent data (Kojevnikov, Marmer and Song, 2021; Leung, 2022). In addition, we provide a heteroskedasticity and autocorrelation-consistent (HAC) variance estimator that is guaranteed to be conservative, enabling researchers to construct Wald-style confidence intervals with valid coverage in large samples.
@@
We explore several extensions of the method, including evaluating the identification assumptions using placebo tests and allowing for diffusion or dependence in treatment assignment. We also discuss the implications of misspecifying the MSM and having measurement error in the proximity metric. We test the method’s performance via Monte Carlo simulations and demonstrate its application through two empirical studies. The first study examines whether wind turbine proposals generate spillover effects in space on political behavior among Canadian voters. The second investigates whether exposure to second-hand smoke through social connections can impact health outcomes in the Framingham Heart Study.
This paper contributes to several strands of the causal inference literature. The definition of the AMR extends the concept of marginalized causal effects (Hudgens and Halloran, 2008; Sävje, Aronow and Hudgens, 2021; Hu, Li and Wager, 2022; Wang et al., 2025) to longitudinal data. We clarify its connection to MSMs and show how it can capture not only direct effects (Papadogeorgou et al., 2020), but also a wide range of spillover effects. Compared to alternative approaches—particularly those based on exposure mappings (Ogburn et al., 2020; Jiang, Wallace and Thompson, 2023; Jetsupphasuk, Li and Hudgens, 2025)—our method requires only minimal knowledge of the interference structure and remains valid even when the effects from other units are complex and heterogeneous. It is also robust to dependence in confounders and contagion in outcomes and avoids the additional step to compute exposure probabilities (Aronow and Samii, 2017; Forastiere, Airoldi and Mealli, 2021).
The estimators we propose mirror the form of classical methods in longitudinal data analysis when interference is absent (Robins, 1986; Robins, Hernan and Brumback, 2000; Cole and Hernán, 2008) and rely on familiar identification assumptions. They are straightforward to implement for applied researchers across biomedical and social sciences and are accompanied by a complete procedure for statistical inference that we extend to estimates of spillover effects (Kojevnikov, Marmer and Song, 2021; Leung, 2022; Gao and Ding, 2023). By incorporating more information from treatment assignment, our approach complements methods built upon outcome models, which are commonly employed in both network and spatio-temporal analyses (Ogburn, Shpitser and Lee, 2020; Reich et al., 2021).
@@ 2.1 Set Up
We focus on complete longitudinal data from either experiments or observational studies with $N$ units spanning over $T+1$ periods, where $N\gg T$ [^3] We discuss scenarios with a large $T$ in Section 5. Throughout the paper, uppercase letters denote random variables, lowercase letters represent their realizations, and boldface indicates vectors or matrices. Subscripts refer to specific units, while superscripts in parentheses indicate specific time periods. For each unit $i\in\mathcal{N}=\{1,2,\dots,N\}$ in period $t\in\{0,1,2,\dots,T\}$, we observe the treatment status $A_{i}^{(t)}\in\{0,1\}$ and a vector of time-varying confounders $\mathbf{L}_{i}^{(t)}$, which can also include time-invariant variables. For simplicity, we assume that $A_{i}^{(0)}=0$ for all $i$. The outcome $Y_{i}$ is measured in the final period $T$. We assume that units in the data are embedded in a fixed social network or geographic space, $\mathcal{G}_{N}$, with a time-invariant proximity metric $d_{ij}$ defined for every pair of units $i$ and $j$.
@@
[…]
Notes: The DAG depicts two units, $\{i,j\}$, observed over two periods, $\{T-1,T\}$. Variables are represented by circles and causal paths by arrows. White circles denote treatment, gray circles represent time-varying confounders, and black circles correspond to outcomes. Red arrows illustrate the effect of a unit’s own treatment history, while blue arrows capture spillover effects due to interference. Black arrows indicate relationships influencing treatment assignment, and gray arrows represent other potential dependencies between variables under the assumptions outlined in Section 3 (e.g., causal mechanisms).
@@
$Y_{i}=Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right),\mathbf{L}_{i}^{(t)}=\mathbf{L}_{i}^{(t)}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right),$ (1)
@@ 2.2 Estimands
Our goal is to understand how the treatment history of a unit influences its own outcome or the outcome of its neighbors in $\mathcal{G}_{N}$. Let $\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}$ represent the remaining part in $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ after removing unit $j$’s treatment history between periods $1$ and $T$. Then, we use $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=Y_{i}\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ to denote the potential outcome of unit $i$ when the treatment history of unit $j$ is fixed at $\mathbf{a}^{(1:T)}$ but all other units’ treatment histories are assigned at random. To construct meaningful estimands, we take the expectation of $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ over $\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}$ and other random variables that may affect $Y_{i}$, obtaining the marginalized potential outcome for unit $i$ when $j$’s treatment history is fixed:
$Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\coloneqq{\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right].\vspace{-0.5em}$ (2)
This quantity can be defined for every pair of $i$ and $j$, including cases where $i=j$.
@@
Next, for each unit $j$, we aggregate $Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ over a pre-specified set of $j$’s neighbors denoted as $\Omega_{j}(d)$, where $d$ is a specific value of the proximity metric. This leads to a linear mapping $\mu\left(\cdot;\Omega_{j}(d)\right)$ from the non-random vector $\big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}}$ to a real number:
@@
$\mu\left(\big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)$ $\coloneqq\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)}{|\Omega_{j}(d)|},$ (3)
@@
where $\mathbf{1}\{\cdot\}$ is the indicator function and $|\Omega_{j}(d)|=\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}$ represents the number of units in $\Omega_{j}(d)$. The form of $\Omega_{j}(d)$ can be chosen by the researcher based on the type of spillover effects under study. In social networks, a natural choice is the $d$th-degree neighbors of each $j$: $\Omega_{j}(d)\coloneqq\{i\in\mathcal{N}:d_{ij}=d\}$. In space, one option is the “circle average” introduced by Wang et al. (2025), where $d$ stands for any distance value and $\Omega_{i}(d)$ represents a “doughnut:” $\{i\in\mathcal{N}:d-\kappa<d_{ij}\leq d\}$, with $\kappa$ being a fixed constant. In either case, the quantity defined in (3) captures the expected response of unit $j$’s neighbors whose proximity to $i$ is (approximately) $d$ to $i$’s treatment history. With a slight abuse of terminology, we will refer to units belonging to $\Omega_{j}(d)$ as $j$’s $d$th-degree neighbors henceforth. When $d=0$, $\Omega_{j}(d)$ only includes unit $j$ itself, and the quantity boils down to $Y_{j;j}\left(\mathbf{a}^{(1:T)}\right)$, unit $j$’s expected outcome given its own treatment history.
@@
In the top-left plot of Figure 2, all the second-degree neighbors ($d=2$) of unit $42$ are marked in red. In the top-right plot, a specific unit $j$ is highlighted in red, while red circles indicate the range of $\Omega_{j}(d)$ (in the form of doughnuts with $\kappa=1$) for $d\in\{0,1,2,3\}$. Both $d$ and $\Omega_{j}(d)$ can be adapted to specific contexts. For example, $d$ may take discrete values indicating whether two units are from the same street or district, or represent traffic accessibility or even cultural proximity between units. We discuss potential impacts of measurement error of $d$ in Section 5 and summarize choices of $d$ and $\Omega_{j}(d)$ in the Supplementary Material. When the context is clear, we represent the value of $\mu(\cdot;\Omega_{j}(d))$ with $\mu_{j}(\cdot;d)$.
@@
It is worth noting that the definition of $\Omega_{j}(d)$ does not require any knowledge of the interference structure—specifically, the functional form of $Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)$. It remains well-defined as long as the proximity metric exists. Instead of explicitly studying how a unit’s outcome is influenced by the others—typically formalized through an exposure mapping—our framework examines the effects of a unit’s treatment history on a predefined set of neighbors. We do not assume that any $\Omega_{j}(d)$ or their collection captures all the influences generated by a unit. The form of $\Omega_{j}(d)$ is flexible and not unique. [^5] But it does impact statistical inference. We discuss this issue in Section 4. Overlap of $\Omega_{j}(d)$ across the units is permitted, hence each unit may contribute to multiple $\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)$ values if it is influenced by more than one neighbor.
@@
Finally, by taking the average of $\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)$ across all units in the sample, we obtain the average marginalized response (AMR):
$\mu\left(\mathbf{a}^{(1:T)};d\right)\coloneqq\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\Big{\{}Y_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right).\vspace{-0.5em}$ (4)
@@
Given a pre-specified range $\mathcal{D}$, the AMR can be defined for each $d\in\mathcal{D}$ and enables researchers to assess how spillover effects vary with proximity without making assumptions about their magnitude. For instance, we could set $\mathcal{D}=\{0,1,2,3,4\}$ if the goal is to test whether spillover effects only extend to one’s third-degree neighbors. Moreover, the AMR at any $d$ depends only on the treatment history $\mathbf{a}^{(1:T)}$, as other factors affecting the outcome have been marginalized over. Thus, when $T$ is not too large, the relationship between any AMR and the corresponding treatment history can be represented by a saturated marginal structural model (MSM):
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}(d).\vspace{-1em}$ (5)
@@
Here, $\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)$ consists of indicators for the treatment status in each period and their interactions, and $\boldsymbol{\beta}(d)$ is a vector of parameters specific to $d$. When $d=0$, Equation (5) resembles a conventional MSM, where the parameters describe the direct effects of a unit’s treatment histories on its own expected outcome. When $d>0$, $\boldsymbol{\beta}(d)$ captures spillover effects those treatment histories generate on the expected outcome of the unit’s $d$th-degree neighbors. [^6] As in a conventional MSM, these quantities include the effects mediated through the time-varying confounders, represented by gray arrows in Figure 1. $\{\boldsymbol{\beta}(d)\}_{d\in\mathcal{D}}$ thus provides a comprehensive view of both direct and spillover effects in the sample.
@@
Similar to other marginalized estimands in the literature (Hudgens and Halloran, 2008; Sävje, Aronow and Hudgens, 2021; Hu, Li and Wager, 2022; Wang et al., 2025), these causal quantities reflect the effect of manipulating a single unit’s treatment history under the current assignment policy, rather than contrasting hypothetical scenarios where all units receive specific treatments. They are inherently sample-specific (though we omit $\mathcal{N}$ in the subscript for simplicity) and “descriptive” of the status quo, not “prescriptive” for identifying the optimal policy (Kennedy, 2019). As such, $\boldsymbol{\beta}(d)$ does not represent structural parameters invariant to changes in the assignment policy. Nevertheless, it remains meaningful under the prevailing policy, capturing all causal pathways through which treatment influences outcomes, including direct effects, spillover effects, changes in time-varying confounders, and shifts in treatment distributions over time.
@@
$AMR$ $\mu\left((0,0);d\right)$ $\mu\left((0,1);d\right)$ $\mu\left((1,0);d\right)$ $\mu\left((1,1);d\right)$

### v5

@@ Abstract
In longitudinal studies where units are embedded in space or networks, interference may arise, meaning that a unit’s outcome can depend on treatment histories of others. The presence of interference poses significant challenges for causal inference, particularly when the interference structure—how a unit’s outcome responds to others’ influences—is complex, heterogeneous, and unknown. This paper develops a general framework for identifying and estimating both direct and spillover effects of treatment histories under minimal assumptions about the interference structure. We introduce a class of causal estimands that capture the effects of treatment histories at any specified proximity level and show that they can be represented by a modified marginal structural model. Under sequential exchangeability, these estimands are identifiable and can be estimated using inverse probability weighting. We derive conditions for consistency and asymptotic normality of the estimators and provide procedures for constructing asymptotically conservative confidence intervals. The method’s utility is demonstrated through applications in both social science and biomedical settings.
@@ 1 Introduction
Empirical researchers are often interested in estimating the causal effects of treatment histories in longitudinal data. For example, public health researchers may seek to determine whether a history of smoking increases the likelihood of developing cardiovascular diseases, while social scientists want to understand how past experiences with government programs shape voting behavior. Under the standard assumption of sequential exchangeability—in either experimental or observational studies—these effects can be represented by a marginal structural model (MSM) and consistently estimated through inverse probability weighting (IPW) (Robins, Hernán and Brumback, 2000; Hernan and Robins, 2025).
@@
However, this classic approach relies on the stable unit treatment value assumption (SUTVA), which rules out the possibility that the outcome of a unit is affected by treatment histories of others, a phenomenon known as spillover or interference (Cox, 1959). In contexts where units are embedded in geographic space or social networks, interference is likely present, and researchers may lack knowledge of the interference structure—the relationship between a unit’s outcome and the treatments of others. In reality, this relationship can be complex and heterogeneous across units. For example, some individuals’ health may be influenced only by secondhand smoke from immediate contacts in the social network, whereas others who play a more active role in the community may also be affected by the smoking behavior of individuals to whom they are not directly connected. In such settings, which we refer to as “unknown interference,” two key questions arise: First, what can still be learned using the classic approach? Second, how can we identify and estimate spillover effects generated by one unit’s treatment on the outcome of others?
In this paper, we present a general framework to address these questions. We introduce a novel estimand, the average marginalized response (AMR), which is defined as the average, across all units in the sample, of the expected outcomes of their neighbors at a proximity level of $d\geq 0$, under a fixed treatment history for the unit. The AMR remains well-defined for any user-specified proximity metric, such as spatial distance, network path length, or shared group membership, even when the interference structure is unknown. [^3] The proximity metric is used solely to index spillover effects and does not impose any restrictions on the form or complexity of the interference structure. When $d=0$, the AMR reduces to the expected outcome of each unit under a fixed treatment history of its own, averaged over the sample. For any $d$, the AMR is a function of treatment history alone, thus can be represented by a MSM, where linear combinations of the parameters describe the causal effect of specific treatment histories on the expected outcome of the unit itself (when $d=0$) or its neighbors at the proximity level of $d$ (when $d>0$) under the prevailing treatment assignment policy.
We then show that both the AMR and the associated MSM parameters can be identified under the standard assumption of sequential exchangeability. This identification result enables estimation of MSM parameters using IPW estimators, such as weighted least squares (WLS), where we regress the average outcome of a unit’s neighbors at the proximity level of $d$ on indicators of its own treatment history, weighted by the inverse probability of that history being observed. When $d=0$, the estimator coincides with the classic approach, so its results can be interpreted as estimated effects for the AMR at 0. Estimates for $d>0$ allow researchers to examine how the effects of different treatment histories extend to neighbors at varying levels of proximity.
@@
We prove that this estimator exhibits desirable large-sample properties—specifically, consistency and asymptotic normality—when either the influence of units’ treatments decays sufficiently quickly over proximity or neighborhood sizes grow slowly with the sample size, using recently developed central limit theorems for dependent data (Kojevnikov, Marmer and Song, 2021; Leung, 2022). In addition, we provide a heteroskedasticity and autocorrelation-consistent (HAC) variance estimator that is guaranteed to be conservative, enabling researchers to construct Wald-style confidence intervals with asymptotically valid coverage.
@@
We explore several extensions of the method, including evaluating the identification assumptions using placebo tests and allowing for diffusion or dependence in treatment assignment. We also discuss the implications of misspecifying the MSM and having measurement error in the proximity metric. We test the method’s performance via Monte Carlo simulations and demonstrate its application through two empirical studies. The first study examines whether wind turbine proposals generate spillover effects on political behavior among Canadian voters. The second study investigates whether exposure to second-hand smoke through social connections can impact health outcomes in the Framingham Heart Study.
This paper contributes to several strands of the causal inference literature. Similar to the result in Sävje, Aronow and Hudgens (2021) for cross-sectional settings, we show that the classic approach in longitudinal analysis (Robins, Hernán and Brumback, 2000; Hernan and Robins, 2025) yields estimates with a causal interpretation even under interference. We further extend this approach to estimate spillover effects at various proximity levels. The definition of the AMR generalizes the concept of marginalized causal effects (Hudgens and Halloran, 2008; Hu, Li and Wager, 2022; Papadogeorgou et al., 2020; Wang et al., 2025) to longitudinal settings, and we clarify its connection to MSMs.
Compared to methods based on “exposure mappings” (Ogburn et al., 2020; Jiang, Wallace and Thompson, 2023; Jetsupphasuk, Li and Hudgens, 2025), which assume that the influence of other units follows a known functional form, our approach requires only minimal knowledge of the interference structure and remains valid when this structure differs across units. It is also robust to dependence in confounders and contagion in outcomes and avoids the additional step of computing exposure probabilities (Aronow and Samii, 2017; Forastiere, Airoldi and Mealli, 2021). We demonstrate how the existing inference procedure developed for exposure mappings (Kojevnikov, Marmer and Song, 2021; Leung, 2022; Gao and Ding, 2023) can be applied to estimates of spillover effects in our framework. By incorporating more information from treatment assignment, our approach complements methods built upon outcome models, which are commonly employed in both network and spatio-temporal analyses (Ogburn, Shpitser and Lee, 2020; Reich et al., 2021).
@@ 2.1 Set Up
We focus on complete longitudinal data from either experiments or observational studies with $N$ units spanning over $T+1$ periods, where $N\gg T$. [^4] We discuss scenarios with a large $T$ in Section 5. Throughout the paper, uppercase letters denote random variables, lowercase letters represent their realizations, and boldface indicates vectors or matrices. Subscripts refer to specific units, while superscripts in parentheses indicate specific time periods. For each unit $i\in\mathcal{N}=\{1,2,\dots,N\}$ in period $t\in\{0,1,2,\dots,T\}$, we observe the treatment status $A_{i}^{(t)}\in\{0,1\}$ and a vector of time-varying confounders $\mathbf{L}_{i}^{(t)}$, which can also include time-invariant variables. For simplicity, we assume that $A_{i}^{(0)}=0$ for all $i$. The outcome $Y_{i}$ is measured in the final period $T$. We assume that units in the data are embedded in a fixed social network or geographic space, $\mathcal{G}_{N}$, with a time-invariant proximity metric $d_{ij}$ defined for every pair of units $i$ and $j$.
@@
$Y_{i}$$Y_{j}$$\mathbf{L}_{i}^{(T)}$$\mathbf{L}_{j}^{(T)}$$A_{i}^{(T)}$$A_{i}^{(T-1)}$$A_{j}^{(T)}$$A_{j}^{(T-1)}$
Notes: The DAG depicts two units, $\{i,j\}$, observed over two periods, $\{T-1,T\}$. Variables are represented by circles and causal paths by arrows. White circles denote treatment, gray circles represent time-varying confounders, and black circles correspond to outcomes. Red arrows illustrate the effect of a unit’s own treatment history, while blue arrows capture spillover effects due to interference. Black arrows indicate relationships influencing treatment assignment, and gray arrows represent other potential dependencies between variables under the assumptions outlined in Section 3.
@@
$Y_{i}=Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right),\mathbf{L}_{i}^{(t)}=\mathbf{L}_{i}^{(t)}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right),\vskip-5.0pt$ (1)
@@ 2.2 Estimands
Our goal is to understand how the treatment history of a unit influences its own outcome or the outcome of its neighbors in $\mathcal{G}_{N}$ under the prevailing treatment assignment policy. Let $\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}$ represent the remaining part in $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ after removing unit $j$’s treatment history between periods $1$ and $T$. Then, we use $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)=Y_{i}\left(\mathbf{A}_{j}^{(1:T)}=\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ to denote the potential outcome of unit $i$ when the treatment history of unit $j$ is fixed at $\mathbf{a}^{(1:T)}$ but all other units’ treatment histories are generated independently from the policy.
To construct meaningful estimands, we take the expectation of $Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)$ over $\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}$. This yields the marginalized potential outcome for unit $i$ given $j$’s treatment history:
$\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)\coloneqq{\rm E}\,\left[Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\right].\vskip-5.0pt$ (2)
This quantity can be defined for every pair of $i$ and $j$, including cases where $i=j$. In what follows, $\rm E$$\cdot$] always denotes the expectation over treatment histories conditional on other variables that may affect the outcome, such as pre-treatment covariates and measurement error. In Section A.1 of the Supplementary Material, we show that this definition extends to expectations taken over all variables affecting the outcome, with our method still applying.
@@
Next, for each unit $j$, we aggregate $\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)$ over a pre-specified set of $j$’s neighbors denoted as $\Omega_{j}(d)$, where $d$ is a specific value of the proximity metric. This leads to a linear mapping $\mu\left(\cdot;\Omega_{j}(d)\right)$ from the non-random vector $\big{\{}\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}}$ to a real number:
@@
$\mu\left(\big{\{}\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}};\Omega_{j}(d)\right)$ $\coloneqq\frac{\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)}{|\Omega_{j}(d)|},\vskip-5.0pt$ (3)
@@
where $\mathbf{1}\{\cdot\}$ is the indicator function and $|\Omega_{j}(d)|=\sum_{i=1}^{N}\mathbf{1}\{i\in\Omega_{j}(d)\}$ represents the number of units in $\Omega_{j}(d)$. The form of $\Omega_{j}(d)$ can be chosen by the researcher based on the type of spillover effects under study. In social networks, a natural choice is the $d$th-degree neighbors of each $j$: $\Omega_{j}(d)\coloneqq\{i\in\mathcal{N}:d_{ij}=d\}$. In spatial settings, one option is the “circle average” introduced by Wang et al. (2025), where $d$ stands for any distance value and $\Omega_{i}(d)$ represents a “doughnut”: $\{i\in\mathcal{N}:d-\kappa<d_{ij}\leq d\}$, with $\kappa$ being a fixed constant. In either case, the quantity defined in (3) captures the expected response of unit $j$’s neighbors whose proximity to $i$ is (approximately) $d$ to $i$’s treatment history. With a slight abuse of terminology, we will refer to units belonging to $\Omega_{j}(d)$ as $j$’s $d$th-degree neighbors henceforth. When $d=0$, $\Omega_{j}(d)$ only includes unit $j$ itself, and the quantity reduces to $Y_{j;j}\left(\mathbf{a}^{(1:T)}\right)$, unit $j$’s expected outcome given its own treatment history.
@@
In the top-left plot of Figure 2, all the second-degree neighbors ($d=2$) of unit $42$ are marked in red. In the top-right plot, a specific unit $j$ is highlighted in red, while red circles indicate the range of $\Omega_{j}(d)$ (in the form of doughnuts with $\kappa=1$) for $d\in\{0,1,2,3\}$. Both $d$ and $\Omega_{j}(d)$ can be adapted to specific contexts. For example, $d$ may take discrete values indicating whether two units are from the same street or district, or represent traffic accessibility or even cultural proximity between units. $\Omega_{j}(d)$ may include all units within the proximity range of $d$, or be further restricted to units with specific covariate profiles to capture conditional effects. [^6] These choices do impact statistical inference. We discuss this issue in Section 4. We discuss potential impacts of measurement error of $d$ in Section 5 and summarize choices of $d$ and $\Omega_{j}(d)$ in Section C.1 of the Supplementary Material. When the context is clear, we represent the value of $\mu(\cdot;\Omega_{j}(d))$ with $\mu_{j}(\cdot;d)$.
@@
It is worth noting that the definition of $\Omega_{j}(d)$ does not require any knowledge of the interference structure—specifically, the functional form of $Y_{i}\left(\mathbf{A}_{\mathcal{N}}^{(1:T)}\right)$. It remains well-defined for any proximity metric chosen by the researcher. Instead of explicitly studying how a unit’s outcome is influenced by the others—typically formalized through an exposure mapping—our framework examines the effects of a unit’s treatment history on a predefined set of neighbors. We do not assume that any $\Omega_{j}(d)$ or their collection captures all the influences generated by a unit. Overlap of $\Omega_{j}(d)$ across the units is permitted, hence each unit may contribute to multiple $\mu_{j}\left(\Big{\{}\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right)$ values if it is influenced by more than one neighbor.
@@
Finally, by taking the average of $\mu_{j}\left(\big{\{}\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)\big{\}}_{i\in\mathcal{N}};d\right)$ across all units in the sample, we obtain the average marginalized response (AMR):
$\mu\left(\mathbf{a}^{(1:T)};d\right)\coloneqq\frac{1}{N}\sum_{j=1}^{N}\mu_{j}\left(\Big{\{}\mu_{i;j}\left(\mathbf{a}^{(1:T)}\right)\Big{\}}_{i\in\mathcal{N}};d\right).\vskip-5.0pt$ (4)
@@
Given a pre-specified range $\mathcal{D}$, the AMR can be defined for each $d\in\mathcal{D}$ and enables researchers to assess how spillover effects vary with proximity without making assumptions about their magnitude. For instance, we could set $\mathcal{D}=\{0,1,2,3,4\}$ if the goal is to test whether spillover effects only extend to one’s third-degree neighbors. Moreover, the AMR at any $d$ depends only on the treatment history $\mathbf{a}^{(1:T)}$ and can therefore be represented by a saturated marginal structural model (MSM) when $T$ is not too large:
$\mu\left(\mathbf{a}^{(1:T)};d\right)=\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)^{\prime}\boldsymbol{\beta}(d).\vskip-10.00002pt$ (5)
@@
Here, $\mathbf{m}\left(\mathbf{a}^{(1:T)}\right)$ consists of indicators for the treatment status in each period and their interactions, and $\boldsymbol{\beta}(d)$ is a vector of parameters specific to $d$. When $d=0$, Equation (5) resembles a classic MSM, where the parameters describe the direct effects of a unit’s treatment histories on its own expected outcome. When $d>0$, $\boldsymbol{\beta}(d)$ captures spillover effects generated by treatment histories on the expected outcome of the unit’s $d$th-degree neighbors. [^7] As in a conventional MSM, these quantities include the effects mediated through the time-varying confounders, represented by gray arrows in Figure 1. $\{\boldsymbol{\beta}(d)\}_{d\in\mathcal{D}}$ thus provides a comprehensive view of both direct and spillover effects in the sample.
@@
Similar to other marginalized estimands in the literature (Hudgens and Halloran, 2008; Sävje, Aronow and Hudgens, 2021; Hu, Li and Wager, 2022; Wang et al., 2025), these causal quantities reflect the effect of manipulating a single unit’s treatment history under the current treatment assignment policy, rather than contrasting hypothetical scenarios where all units receive specific treatments. They are inherently sample-specific (though we omit $\mathcal{N}$ in the subscript for simplicity) and “descriptive” of the status quo, not “prescriptive” for identifying optimal policies (Kennedy, 2019). As such, $\boldsymbol{\beta}(d)$ does not represent parameters that are invariant to policy changes. Nevertheless, it remains meaningful under the prevailing policy, capturing all causal pathways through which treatment influences outcomes, including direct effects, spillover effects, changes in time-varying confounders, and shifts in treatment distributions over time. In Section A.2 of the Supplementary Material, we illustrate how the AMRs can inform the direction of policy changes that improve the average expected outcome, also known as welfare in the literature (Hu, Li and Wager, 2022; Viviano, 2020).
@@
AMRAMR $\mu\left((0,0);d\right)$ $\mu\left((0,1);d\right)$ $\mu\left((1,0);d\right)$ $\mu\left((1,1);d\right)$